# Identifying vegetation patterns for a qualitative assessment of land degradation using a cellular automata model and satellite imagery

### v2

@@ Abstract
We aim to identify the spatial distribution of vegetation and its growth dynamics with the purpose of obtaining a qualitative assessment of vegetation health and land degradation. To do so, we compare a statistical model of vegetation growth and real-world Earth observation data. Specifically, we analyze a stochastic cellular automata model and data obtained from satellite images, namely using the Normalized Difference Vegetation Index (NDVI) and the Leaf Area Index (LAI). In the experimental data, we look for areas where vegetation is broken into small patches and qualitatively compare it to the percolating, fragmented, and degraded states that appear in the cellular automata model. We model the periodic effect of seasons, finding numerical evidence of a periodic fragmentation and recovery phenomenology if the model parameters are sufficiently close to the model’s percolation transition. We qualitatively recognize these effects in real-world vegetation images and consider them a signal of increased environmental stress and vulnerability. Finally, we show an estimation of the environmental stress in land images by considering both the vegetation density and its clusterization.
@@ I Introduction
Monitoring large land areas is fundamental to understand the progression of these environmental changes. To this aim, numerous studies reynolds2007global ; sala2000global ; archer2001trees have explored the relationships among drylands, vegetation patterns, and external stresses, using a number of models and data sources. These include remote sensing jensen2009remote , which uses satellite or airborne sensors to gather data on vegetation cover and health; geographic information systems (GIS) goodchild1991geographic for analyzing spatial patterns and changes in vegetation; ecohydrological modeling tague2004rhessys to examine the interaction between vegetation and water resources; agent-based modeling (ABM) grimm2006standard to simulate individual-level behavior and its impact on ecosystems; and species distribution modeling (SDM) elith2006novel to predict plant species distribution under different environmental conditions. A statistical physics approach, alongside these methods, also contributes to understanding vegetation distribution and dynamics, as well as the role of disorder in the vegetation processes. Particularly, vegetation patchiness is a valuable tool for assessing degradation risk and detecting early warning signals associated with it levick2013 ; dahlin2013 ; asner2013forest . Disorder enters the vegetation dynamics through environmental fluctuations, stochastic processes such as weather patterns, disturbances (e.g., wildfires, human activities), and spatial heterogeneity in soil composition or topography. The presence of disorder affects both the existence and the way the transitions occurs, making them more gradual or continuous villa2015eluding ; martin2014quenched ; vazquez2011temporal . This, in turn, may affect the environmental resilience.
@@
Our objective in this work is to describe the structure of vegetation cover through a numerical model of vegetation growth dynamics, and to conduct an analysis of satellite image data to recognize patterns of vegetation patchiness that are signs of land degradation in progress. We use a Stochastic Cellular Automaton (SCA) model kefi2007spatial ; kefi2010bistability ; kefi2007local ; kefi2014early , which incorporates events such as vegetation mortality, survival and propagation. This SCA model is born out of individual plant lifecycle considerations (e.g. propagation of seeds, new plant colonization, plant competition, etc). These effects are coarse-grained into stochastic state transitions of macroscopic cells which are arranged in a grid, covering large land areas. The state changes in the cells are able to generate patterns observed in arid ecosystems, such as gaps, stripes, and labyrinth-like structures kefi2007spatial ; kefi2007local . It also presents two phase transitions as the environmental stress (or vegetation mortality) changes: a desertification transition where vegetation disappears, and a percolation transition where the vegetation breaks down into clusters and which acts as a precursor to the desertification transition corrado2014early ; manor2008facilitation ; von2010periodic .
@@
The work is structured as follows: Section II provides an overview of the theoretical and numerical methods employed, including the model used for simulations and the incorporation of seasonal effects. In Section III, the analysis is applied to recent satellite images of a selection of European countries. Section IV connects the SCA simulated process and the observed land data. Finally, our conclusions are outlined in Section V.
@@ II.1 Numerical simulations of vegetation dynamics
In this section, we investigate the model for different values of the mortality rates $m$. The inset in Fig. 1(a) shows the evolution of the living cell density $\rho_{+}(t)$, for several values of $m$. After a short transient (not shown), the system fluctuates around an equilibrium value $\rho_{+}(m)$. With the increase of external stress $m$, the average density of vegetated cells decreases, until a continuous phase transition occurs at $m_{c}=0.169$ (see the main panel of Fig. 1(a)) corresponding to a desertification transition. Additionally, we determine the size of the largest cluster of vegetated cells $\mathcal{C}_{+}(m)$. As $m$ increases, at finite sizes $\mathcal{C}_{+}$ has a sharp crossover that corresponds to a percolation transition which acts as a precursor to desertification corrado2014early ; kefi2007spatial ; kefi2014early . Considering different system sizes allows to find the percolation transition point at $m_{Per}=0.113$ as the crossing point of the curves at different $L$. See the Supplementary Material (SM) Sec. I for additional results on these transitions, particularly the critical exponents.
Thus, two distinct transitions are involved in land degradation processes corrado2014early ; corrado2015desertification ; corrado2015critical ; bengochea2022habitat . The first transition, a percolation transition in the living vegetation clusters, serves as an early warning sign for the second transition, which is associated with desertification as the fraction of living vegetation goes to zero. Both of these transitions have been observed to occur in real ecosystems von2010periodic . The bottom panels of Fig. 1 shows snapshots of the model at different stages: healthy and percolating (panel b), vulnerable and broken into small clusters (panel c), and degraded (panel d).

### v3

@@ Abstract
We aim to identify the spatial distribution of vegetation and its growth dynamics with the purpose of obtaining a qualitative assessment of vegetation characteristics tied to its condition, productivity and health, and to land degradation. To do so, we compare a statistical model of vegetation growth and land surface imagery derived vegetation indices. Specifically, we analyze a stochastic cellular automata model and data obtained from satellite images, namely using the Normalized Difference Vegetation Index (NDVI) and the Leaf Area Index (LAI). In the experimental data, we look for areas where vegetation is broken into small patches and qualitatively compare it to the percolating, fragmented, and degraded states that appear in the cellular automata model. We model the periodic effect of seasons, finding numerical evidence of a periodic fragmentation and recovery phenomenology if the model parameters are sufficiently close to the model’s percolation transition. We qualitatively recognize these effects in real-world vegetation images and consider them a signal of increased environmental stress and vulnerability. Finally, we show an estimation of the environmental stress in land images by considering both the vegetation density and its clusterization.
@@ I Introduction
Monitoring large land areas is fundamental to understand the progression of these environmental changes. To this aim, numerous studies reynolds2007global ; sala2000global ; archer2001trees have explored the relationships among drylands, vegetation patterns, and external stresses, using a number of models and data sources. These include remote sensing jensen2009remote , which uses satellite or airborne sensors to gather data on vegetation cover and health; geographic information systems goodchild1991geographic for analyzing spatial patterns and changes in vegetation; ecohydrological modeling tague2004rhessys to examine the interaction between vegetation and water resources; agent-based modeling grimm2006standard to simulate individual-level behavior and its impact on ecosystems; and species distribution modeling elith2006novel to predict plant species distribution under different environmental conditions. A statistical physics approach, alongside these methods, also contributes to understanding vegetation distribution and dynamics, as well as the role of disorder in the vegetation processes. Particularly, vegetation patchiness is a valuable tool for assessing degradation risk and detecting early warning signals associated with it levick2013 ; dahlin2013 ; asner2013forest . Disorder enters the vegetation dynamics through environmental fluctuations, stochastic processes such as weather patterns, disturbances (e.g., wildfires, human activities), and spatial heterogeneity in soil composition or topography. The presence of disorder affects both the existence and the way the transitions occurs, making them more gradual or continuous villa2015eluding ; martin2014quenched ; vazquez2011temporal . This, in turn, may affect the environmental resilience.
@@
Our objective in this work is to describe the structure of vegetation cover through a numerical model of vegetation growth dynamics, and to conduct an analysis of satellite image data to recognize patterns of vegetation patchiness that are signs of land degradation in progress. We use a Stochastic Cellular Automaton (SCA) model kefi2007spatial ; kefi2007local ; kefi2010bistability ; kefi2014early , which incorporates events such as vegetation mortality, survival and propagation. This SCA model is born out of individual plant lifecycle considerations (e.g. propagation of seeds, new plant colonization, plant competition, etc). These effects are coarse-grained into stochastic state transitions of macroscopic cells which are arranged in a grid, covering large land areas. The state changes in the cells are able to generate patterns observed in arid ecosystems, such as gaps, stripes, and labyrinth-like structures kefi2007spatial ; kefi2007local . It also presents two phase transitions as the environmental stress (or vegetation mortality) changes: a desertification transition where vegetation disappears, and a percolation transition where the vegetation breaks down into clusters and which acts as a precursor to the desertification transition corrado2014early ; manor2008facilitation ; von2010periodic .
@@
The work is structured as follows: Sec. II provides an overview of the theoretical and numerical methods employed, including the model used for simulations and the incorporation of seasonal effects. In Sec. III, the analysis is applied to recent satellite images of a selection of European countries. Section IV connects the SCA simulated process and the observed land data. Finally, our conclusions are outlined in Sec. V.
@@ II.1 Numerical simulations of vegetation dynamics
In this section, we investigate the model for different values of the mortality rates $m$. The inset in Fig. 1(a) shows the evolution of the living cell density $\rho_{+}(t)$, for several values of $m$. After a short transient (not shown), the system fluctuates around an equilibrium value $\rho_{+}(m)$. With the increase of external stress $m$, the average density of vegetated cells decreases, until a continuous phase transition occurs at $m_{c}=0.169$ (see the main panel of Fig. 1(a)) corresponding to a desertification transition. Additionally, we determine the size of the largest cluster of vegetated cells $\mathcal{C}_{+}(m)$. As $m$ increases, at finite sizes $\mathcal{C}_{+}$ has a sharp crossover that corresponds to a percolation transition which acts as a precursor to desertification corrado2014early ; kefi2007spatial ; kefi2014early . Considering different system sizes allows to find the percolation transition point at $m_{Per}=0.113$ as the crossing point of the curves at different $L$. See Sec. I of Supplemental Material for information about the transitions.
Thus, two distinct transitions are involved in land degradation processes corrado2014early ; corrado2015desertification ; corrado2015critical ; bengochea2022habitat . The first transition, a percolation transition in the living vegetation clusters, serves as an early warning sign for the second transition, which is associated with desertification as the fraction of living vegetation goes to zero. Both of these transitions have been observed to occur in ecosystems von2010periodic . The bottom panels of Fig. 1 shows snapshots of the model at different stages: healthy and percolating (panel b), vulnerable and broken into small clusters (panel c), and degraded (panel d).